OpenStudy (anonymous):

On a single toss of a die, what is the probability of obtaining the following? An even number & a number less than 5.

OpenStudy (m):

5/6

OpenStudy (anonymous):

Even # : 3/6 Less than 5: 4/6

OpenStudy (m):

probability of getting number < 5 = 4/6 getting even = 3/6 but 2 and 4 already counted so 5/6

OpenStudy (anonymous):

Both Combined: 2/6